OpenStudy (anonymous):

Find the logarithm. log_4(1/64)=?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

let \(y=log_4(1/64)\) Then you can do \(4^y=1/64\) \(1/64=64^{-1}\) and \(64=4^3\) so \(1/64=4^{-3}\) \[4^y=4^{-3}\] so \(y=-3\)

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Daniel I have another for you to look at please... :)